OpenStudy (anonymous):

\[Q=Q_c+Q_p\] \[ =e^{-10t}(c_1cos20t+c_2sin20t)+\frac{3}{125}\] \[Q(0)=0\] \[Q'(0)=0\]

OpenStudy (anonymous):

how do I find c_1 and c_2?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

I'm almost done solving the problem, I just don't know how to use the initial charge and current (both are zero)

zepdrix (zepdrix):

\[\large Q(0)=0 \quad \rightarrow \quad 0=e^{-10\cdot0}(c_1\cos20\cdot0+c_2\sin20\cdot0)+\frac{3}{125}\] Hmm yah both are zero :D Interesting. You shouldn't any problem though, since cosine 0 will give you something useful. As long as you understand where to plug in the 0's :D

OpenStudy (anonymous):

cosine of zero is 1, and e^0 is one. 1=-3/125?

OpenStudy (anonymous):

oooops

zepdrix (zepdrix):

for c1? :D yah looks good! c2 will be a bit more work.. since you need to find Q' :C

OpenStudy (anonymous):

I meant c_1

OpenStudy (anonymous):

Q'c=0

OpenStudy (anonymous):

I mean Q'(0)=0

zepdrix (zepdrix):

Yes so we need our equation in terms of Q' so we can plug in the 0's

zepdrix (zepdrix):

\[\large Q(t)=e^{-10t}(c_1\cos20t+c_2\sin20t)+\frac{3}{125}\]\[\large Q'(t)=?\]

zepdrix (zepdrix):

Bunch of product rule shinanigans :O

OpenStudy (anonymous):

\[Q'=-10e^{-10t}(c_1cos20t+c_2sin20t)+e^{-10t}(-20c_1sin20t+20c_2cos20t)\] \[\]

zepdrix (zepdrix):

So now with the derivative, you might notice that you'll actually have 2 terms to worry about when you plug in zeros at this point - 2 terms that give you non-zero since the product rules produced a couple cosine terms. So just a little bit more work to find c2! :)

OpenStudy (anonymous):

\[Q'(0)=-10e^{-10(0)}(c_1)+e^{-10(0)}20c_2\] \[Q'(0)=c_1+20c_2=0\]

OpenStudy (anonymous):

do I have to take one more derivative

zepdrix (zepdrix):

no c:

zepdrix (zepdrix):

If there was a third unknown constant, and a third initial condition, then yah you might have to do that. But that would only show up in a higher order DE :D

OpenStudy (anonymous):

\[c_2=\frac{-c_1}{20}\]

zepdrix (zepdrix):

And we already know what c1 is right? :D plug 'er in!

OpenStudy (anonymous):

\[c_2=\frac{3}{2500}\]

zepdrix (zepdrix):

From here,\[Q'(0)=-10e^{-10(0)}(c_1)+e^{-10(0)}20c_2\]Shouldn't this give us,\[\large c_2=\frac{c_1}{2}\]

OpenStudy (anonymous):

oh I see I neglected the -10. makes sense =D THanks sooo much!!!

zepdrix (zepdrix):

heh :D